OpenStudy (anonymous):

Really Tricky Separable Equation dy/dx=e^(x+y)/(y-1)

OpenStudy (anonymous):

\[\frac{ dy }{ dx }=\frac{ e ^{x+y} }{ y-1 }\]

zepdrix (zepdrix):

Recall your rule for exponents:\[\Large\bf\sf e^{x+y}\quad=\quad e^x\cdot e^y\]

zepdrix (zepdrix):

Do you think you can separate it from there? :o

OpenStudy (anonymous):

ln'ing both sides?

OpenStudy (anonymous):

what i did was i divided both sides by \[\frac{ e ^{y} }{ y-1 }\] but it gets messy

zepdrix (zepdrix):

Yah that seems to be a good approach.\[\Large\bf\sf \frac{dy}{dx}\quad=\quad \frac{e^x\cdot e^y}{y+1}\qquad\to\qquad (y+1)e^{-y}\;dy\quad=\quad e^x\;dx\]

OpenStudy (anonymous):

from there integration by parts?

zepdrix (zepdrix):

Hmmm ya looks like Parts on the left side.

OpenStudy (anonymous):

give me 10000 secs please

zepdrix (zepdrix):

lol :3

zepdrix (zepdrix):

Oh it was (y-1) :P sorry bout the typo before.

OpenStudy (anonymous):

(-y-1)e^-y + e^-y

zepdrix (zepdrix):

Hmm I think you should be getting:\[\Large\bf\sf -(y-1)e^{-y}-e^{-y}\]

OpenStudy (anonymous):

is that with the y+1 error?

OpenStudy (anonymous):

sorry

zepdrix (zepdrix):

No I used (y-1) for the factor :o

zepdrix (zepdrix):

u=(y-1) dv=e^{-y}

OpenStudy (anonymous):

wait also shouldn't it be respect to x? not y??

OpenStudy (anonymous):

im lost im sorry

zepdrix (zepdrix):

We're integrating with respect to y, that's what the differential `dy` is telling us.

zepdrix (zepdrix):

\[\Large\bf\sf \int\limits(y+1)e^{-y}\;dy\quad=\quad \int\limits e^x\;dx\]So this is where we're at?

zepdrix (zepdrix):

So the right side gives us:\[\Large\bf\sf \int\limits\limits(y+1)e^{-y}\;dy\quad=\quad e^x+c\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

should it be y-1 again?

OpenStudy (anonymous):

gomenasai

zepdrix (zepdrix):

Ohh I copy pasted from the start, my bad lolol

OpenStudy (anonymous):

thats cool i still get it

zepdrix (zepdrix):

\[\Large\bf\sf \int\limits\limits\limits(y-1)e^{-y}\;dy\quad=\quad e^x+c\] \[\Large\bf\sf u=(y-1), \qquad\qquad\qquad dv=e^{-y}\;dy\]\[\Large\bf\sf du=\;dy,\qquad\qquad\qquad\qquad v=-e^{-y}\]

zepdrix (zepdrix):

\[\Large\bf\sf uv-\int\limits v\;du \qquad\to\qquad (y-1)\left(-e^{-y}\right)-\int\limits \left(-e^{-y}\right)dy\]

OpenStudy (anonymous):

1sec im doing it on paper

OpenStudy (anonymous):

oaky you're right

zepdrix (zepdrix):

Simplify things down a tad: \[\Large\bf\sf -(y-1)e^{-y}-e^{-y}\quad=\quad -ye^{-y}\]So we have:\[\Large\bf\sf -ye^{-y}\quad=\quad e^{x}+c\] \[\Large\bf\sf y e^{-y}\quad=\quad C-e^x\]

zepdrix (zepdrix):

Which ... unfortunately doesn't have an explicit form that we can get it into :(

OpenStudy (anonymous):

:)))))))))))))))))))) so that must be the end of the problem then

OpenStudy (anonymous):

as a differential equation professor would you be acceptable of that answer?

zepdrix (zepdrix):

Yes unless you were given initial data that you can plug in to solve for c :)

zepdrix (zepdrix):

For Diff Eq? Yah that's a fine answer. There are some weird things that would allow you to write it terms of x, but not anything that you'll learn in DE.

OpenStudy (anonymous):

this is just homework, but i worry, that my prof will be dissatisfied and give me an F

OpenStudy (anonymous):

bless, thank you very much for you're help!

zepdrix (zepdrix):

Np c: hopefully professor is a nice guy lol